The Table below shows the total demand for cable TV subscriptions for a monopoly. Assume that the monopolist incurs an annual fixed cost of $100,000 and that the marginal cost of providing an additional subscription is always $100.

Quantity

Price (per year)

0

$400

2,000

$350

4,000

$300

6,000

$250

8,000

$200

10,000

$150

12,000

$100

14,000

$50

16,000

$0

    

  1. What is the profit maximising level of quantity and price? What is the profit at this level of output? Explain your answers using the concept of marginal revenue and marginal cost.                                                                                                                 (4 marks)
  2. The following table shows a game played between two players, A and B. The payoffs in the table are shown as (Payoff to A, Payoff to B).

B

Right

Left

A

Up

(2, 2)

(3, 3)

Down

(1, 1)

(4, 0)

Critically analyse the following (remember to justify your answers):

  1. Is Up-Right a Nash equilibrium?                                                       (1 mark)
  2. Is Up-Left a Nash equilibrium?                                                          (1 mark)
  3. Is Down-Right a Nash equilibrium?                                                   (1 mark)
  4. Is Down-Left a Nash equilibrium?                                                     (1 mark)
Answer #1

question (6) MR mc quantity Price TR $400 O $100 2,000 350 700,000 350 100 4,000 300 12,00,000 250 100 6,000 250 15,00,000 15B If P-2 chooses last Then P-1 will choose Down, Profit = TR-TC = 15,00,000- 7,00,000 8,00,000 firm Produce. The output Corre0 (10) cef- Left is a NE? If Pol, chooses up, Then P-2 will choose Left and If f2 chooses Left, Then fI will choose down. Prosorry for inconvenience please take player-1 as player-A and player-2 as player-B !